The problem is typed just as I read it in my daughters homework. What am I missing?
Dottie found 13 leaves in total, in 7 different sizes. Assuming she found the same number of leaves of each size, how many leaves did she find in each size?

Answer by TimothyLamb(4379)

leaf counts must be whole numbers (1, 2, 3, etc.), not fractions (1/2, 3/4, etc.)
---
the problem says: she found leaves in 7 different sizes
---
the problem says: she found the same number of leaves of each size
---
so she had to have found 7*1 or 7*2, or 7*3, etc. number of leaves in total
---
7*1 = 7
7*2 = 14
7*3 = 21
---
13 doesn't match any of the above, so the problem statement contains an error!
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Convert fractions, decimals, and percents:
https://sooeet.com/math/fraction-decimal-percent.php
---
Calculate and graph the linear regression of any data set:
https://sooeet.com/math/linear-regression.php